OpenStudy (anonymous):

x + 4 2x+1 ______ - ______ 2x 6x How to do this?

OpenStudy (ash2326):

Is this your question? \[\frac{x+4}{2x}-\frac{2x+1}{6x}=0\]

OpenStudy (anonymous):

There is no = 0, I just need to simplify it. Thanks!

OpenStudy (ash2326):

Let's take LCM of 2x and 6x \[2x=2\times x\] \[6x=2\times 3 \times x\] LCM= Product of common factors \(\times\) Product of uncommon factors \[common factors= 2 \times x\] uncommon factors=3 so we get \[LCM=2x\times 3=6x\] Now \[\frac{x+4}{2x}-\frac{2x+1}{6x}\] Multiply the numerator of first term by \(\frac{6x}{2x}=3\) and multiply the numerator of second term by\(\frac{6x}{6x}=1\) so we get \[\frac{3\times (x+4)}{6x}-\frac{2x+1}{6x}\] we get \[\frac{3x+12-(2x+1)}{6x}\] We get \[\frac{x+11}{6x}\]

OpenStudy (anonymous):

Wow I get it! Thanks so much!!!!!

OpenStudy (ash2326):

Welcome :D

OpenStudy (anonymous):

What if this problem is multiplying instead of subtracting?

OpenStudy (ash2326):

Then Multiply the numerators separately and multiply the denominators separately

OpenStudy (anonymous):

X+4 2X+1 3X +5 ___ X _____ = ______ Correct?? 2X 6X 12X

OpenStudy (anonymous):

2X + 5 ______ I bet this is right instead 12X

OpenStudy (ash2326):

Cayenne \[(x+4)(2x+1)=x(2x+1)+4(2x+1)=2x^2+2x+8x+1=>2x^2+10x+1\] and \[2x\times 6x= 12x^2\]

OpenStudy (ash2326):

Sorry I made a mistake in the first one \[(x+4)(2x+1)=2x^2+2x+8x+4=2x^2+10x+4\]

OpenStudy (anonymous):

I get x times 2x = 2x squared, but x times 1 - how is that 2x? Thanks

OpenStudy (phi):

X+4 2X+1 2x^2+9X+4 ___ X _____ = ______ 2X 6X 12X^2

OpenStudy (ash2326):

Sorry I made a mistake ,@phi 's answer is correct

OpenStudy (anonymous):

Thanks you are grand! Very helpful...

OpenStudy (ash2326):

Welcome @Cayenne

OpenStudy (anonymous):

If I divided this same equation, I would invert the second fraction and multiply, right?

OpenStudy (ash2326):

Yeah right:D

OpenStudy (phi):

If you divide by a/b invert and multiply by b/a

OpenStudy (anonymous):

Thanks!